OpenStudy (anonymous):

Simplify.[2 × (10 + 5)] - 5 A. 12.5 B. 20 C. 25 D. 120

OpenStudy (janu16):

First, add what is in the parenthesis. (10+5)= ?

OpenStudy (anonymous):

15

OpenStudy (janu16):

Second, fill in 15 in the equation where (10+5) was. [2*15]-5 Third, multiply 2*15. Second, fill in 15 in the equation where (10+5) was. [2*15]-5 Third, multiply 2*15. which is ??

OpenStudy (anonymous):

30

OpenStudy (janu16):

2*15=30 Fourth, fill in 30 where [2*15} was. 30-5 Last, subtract 5 from 30. 30-5=??

OpenStudy (anonymous):

so 25?

OpenStudy (janu16):

yes so C

OpenStudy (janu16):

hope that helped

OpenStudy (anonymous):

thanks!

OpenStudy (skullpatrol):

Good job :-)

OpenStudy (janu16):

yw